In solving the equation (x + 1)(x – 2) = 4, Eric stated that the solution would be
x + 1 = 4 => x = 3
or
(x – 2) = 4 => x = 6
However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

Answer by shalon(2)

(x+1)(x-2)=4
(x^2)-2x+x-2=4 (FOIL)
(x^2)-x-6=0 (move 4 to left side made it general form)
(x+2)(x-3)
x+2=0 or x-3=0
x=-2 x=3
check answer:(x=-2)
(-2+1)(-2-2)=4
-1(-4)=4
4=4 check!
(x=3)
(3+1)(3-2)=4
4(1)=4
4=4 check!